EVA Report – February 1st

EVA #: 19
Author: Scott Dorrington & Steven Hobbs
Start time: 1000
End time: 1230
This EVA was conducted at Candor Chasma. The rovers possessed sufficient charge to allow sufficient time for the crew to penetrate deeply into the Chasma. Total traverse into Candor Chasma was 500 m, with the crew identifying layers of chicken wire gypsum approximately 4 m below the overlying caprock. Brecciated jaspers were also identified, along with layers of purple, pink and green siltstone. Similar siltstones had previously been collected on earlier EVAs and sampled using the VIS/NIR spectrometer instrument, both solid and powdered form.
No fossils were identified from the overlying strata during the latter part of the traverse. Fossilised ripples were identified from an overhanging boulder, suggesting fluid movement too rapid for fossil formation.
Destination: Candor Chasma.
Purpose of EVA: Attempt to identify fossils by exploring deeper into Candor Chasma, in addition to performing a visual characterisation of the exposed strata for future exploration.
Coordinates (use UTM WSG 84): Entrance to Candor Chasma 520779E, 4251127N. Parking at 519868, 4251428. Maximum penetration 520351E, 4251179N
EVA Participants: Steven Hobbs, Scott Dorrington, Alex Tobal
Road(s) and routes per MDRS Map: Exit Hab via main airlock. Take rovers down the driveway to Cow Dung Rd 0110. Take Cow Dung Rd 0110 North to Galileo Rd 1104, turn south past Compass Rock until rover track stops and then walk along the unnamed track towards Candor Chasma, and into Candor Chasma.
Mode of travel: Rover/foot.
Vehicles you will be using (If applicable): Spirit, Opportunity

Supplemental Operations Report – February 4th

Date: 2/4/2024
Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: Perseverance

Power system:
Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 73%; Avg 85.1%.
VDC Last 24 hours: Max 59.5V; Min 42.61; Avg 51.43V.
Generator run time: 4961.8 hours. Serviced on 01/30/2024, oil changed and air filter changed

Propane Readings:
Station Tank: 75%
Director Tank: 71%
Intern Tank: 80%
Generator Tank: 50%

Hab Static Tank – 550 gallons
GreenHab – 195 gallons
Outpost tank – 350 gallons
Science Dome – 0 gallons
Septic Tanks emptied: No

Sojourner rover used: No.
Hours: 196.2
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: Perseverance is not fully charging, inspection will be conducted next week.

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies. Need to be serviced.
Crew Car used and why, where: To Grand Junction by crew
General notes and comments: N/A.

Summary of Internet: All nominal.
EVA suits and radios:
Suits: All nominal
Comms: All nominal
Campus wide inspection, if action taken, what and why: Weather station replaced
Summary of Hab Operations: UV lamp and quartz sleeve in water filter replaced, timer reset. Main heater safety switch replaced. Forehead thermometer in the first aid kit replaced. Toilet sit replaced.
Summary of GreenHab Operations: Night light installer with remote switch. Fan plugged in with remote switch and automated. Supplemental lights plugged in with remote switch and automated.
Summary of SciDome Operations: All nominal
Summary of Observatories Operations: Solar observatory telescope diagonal has two snapped screws. Repairs pending.
Summary of RAM Operations: All nominal.
Summary of Outpost Operations: Intern trailer heater needs to be picked up from Grand Junction. Intern trailer toilet water module assembly replaced.
Summary of Health and Safety Issues: All nominal.

Copyright © The Mars Society. All rights reserved. | Main Site